OpenStudy (anonymous):

a passbook savings account pays a nominal rate of 7% on savings deposits. find the effective annual yield if the interest is compounded semiannually.

OpenStudy (wolf1728):

Annual Rate = (1+.07/2)^2 -1 Annual Rate = 1.071225 -1 Annual Rate = 7.1225%